# Differentially Private Representation Learning via Image Captioning

### v1

@@ 4.1 Downstream Tasks
Zero-shot image classification (V-L) is one of the most widely used methodologies for evaluating vision-language models (Radford et al., 2021). A strong zero-shot performance suggests that the image representation aligns well to text. We perform zero-shot classification using the DP-Cap image encoder and text decoder by evaluating the likelihood of captions of the form “this is a photo of a [label]”. We enumerate over different labels and predict the class that has the highest likelihood; see Section 6.3.1 for full details.
@@ 4.2 Experimental Setup
We present an overview of the experimental setup; refer to Appendix 6 for additional details.
Table 2: Linear probing evaluation on downstream classification. Results for DP-NFNet, TAN, AlexNet and SimCLR are obtained from Yu et al. (2023). For ViP (Yu et al., 2023), we train with the same privacy parameters as for DP-Cap on the deduplicated dataset. More details are given in Appendix 6.1.
@@
Table 3: Performance of DP-Cap on zero-shot classification and compositional understanding (ARO). CLIP’s zero-shot results are obtained from Radford et al. (2021) (base model). For ARO, see Appendix 6.3.2.
@@
We use the ImageNet-1K (Deng et al., 2009a; Russakovsky et al., 2014), CIFAR-10/100 (Krizhevsky et al., 2009), Places-365/205 (Zhou et al., 2014) and iNaturalist-2021 (Van Horn et al., 2021) image classification datasets to assess the performance of learned image representations via full linear probing, few-shot linear probing, and zero-shot prediction. For vision-language tasks, we employ the Visual Genome Attribution (VGA), Visual Genome Relation (VGR), COCO-order (Lin et al., 2015) and Flickr-30k (Plummer et al., 2016) datasets from the ARO benchmark (Yuksekgonul et al., 2022). Finally, we evaluate image captioning using the MS-COCO 2017 (Lin et al., 2015) test set; result is shown in Figure 1(c) and Appendix 7.3.
Model and training. We use a transformer architecture (Vaswani et al., 2017) for both the encoder and the decoder of DP-Cap, where the decoder applies causal cross-attention; see Section 6.1 and Tschannen et al. (2023) for details. For privacy accounting we use Rényi DP composition along with privacy amplification via Poisson subsampling (Mironov et al., 2019), and convert to DP using Balle et al. (2020) through the PyTorch-based Opacus library (Yousefpour et al., 2021), targeting $\delta=1/N$ where $N$ represents the number of training samples. We refer to the non-private counterpart of DP-Cap trained on the same Dedup-LAION-233M dataset as “Cap”.
@@ 4.3 Main Results
Linear probing evaluation (V). We assess the performance of the vision encoder on downstream tasks via linear probing. In Fig 1(a), we compare the performance of DP-Cap and ViP (Yu et al., 2023) on ImageNet-1k few-shot linear probing. DP-Cap significantly improves over ViP, with up to $\times$2.5 better performance across different shots. In addition, we evaluate the full linear probing accuracy of DP-Cap, ViP and other baselines in Table 2. DP-Cap outperforms ViP and other DP models, including TAN (Sander et al., 2023) and DP-NFNet (De et al., 2022), across all tasks. DP-Cap even outperforms non-private AlexNet (Krizhevsky et al., 2012) and except on ImageNet, SimCLR (Chen et al., 2020) (both were trained on ImageNet). We provide additional results for fine-tuning on downstream datasets in Table 10 (App. 7), also showing improvements over competing methods.
@@ 4.4 Ablation Studies
We perform ablation studies on the scaling behavior of DP-Cap with respect to the dataset size, privacy budget and model size. In Appendix 7, we show additional results on image captioning and on the impact of compute budget.
@@
Impact of the privacy budget $\varepsilon$. We also investigate the performance of DP-Cap under lower privacy budgets ($\varepsilon=1$ and $\varepsilon=2$), employing the same batch size of 1.3 million. The outcomes of these experiments are presented in Table 4. As anticipated, the utility of our model does exhibit a decline with decreasing $\varepsilon$. However, the performance degradation is relatively minor for the learned representation, with 10-shot ImageNet performance decreasing from 31.8% ($\varepsilon=8$) to 19.9% ($\varepsilon=1$). More surprisingly, the performance impact on ARO is nearly negligible. It is noteworthy that both models continue to outperform previous state-of-the-art DP models trained with $\varepsilon=8$ (see Figure 1). This phenomenon can be attributed to the relatively small effective noise resulting from the extreme batch size, which for $\varepsilon=1$ remains five times smaller than that used in Yu et al. (2023).
Scaling model size. Scaling up the model size is one of the most effective approaches for training better non-private foundation models (Brown et al., 2020; Bommasani et al., 2021; Touvron et al., 2023). However, conventional wisdom suggests that scaling up model size does not improve utility in DP training since more model parameters will lead to lower signal-to-noise ratio [^1] This is because the added noise has $L_{2}$ norm $\approx\sigma C\sqrt{d}/B$, where $d$ is the number of model parameters, whereas the gradient norm is constrained to $C$ regardless of model size.. To test this hypothesis, we train DP-Cap with different model sizes (Tiny, Small, Base, Large) using the same hyperparameters and evaluate their performance in Table 5,11; see Table 6 for details about different model sizes. We observe consistent improvements when scaling up the model from DP-Cap-Tiny to DP-Cap-Large. Our observation suggests that DP-Cap has strong model scaling behavior even with DP-SGD training.
@@ 5 Discussion and Future Work
Through our study we also identify three open problems in the general direction of DP pre-training of large-scale foundation models that are difficult to handle with existing techniques: 1. While we made notable efficiency improvements to support extreme batch sizes, it remains computationally demanding compared to non-private training. Is it possible to do DP training without the use of extreme batch sizes? 2. Are there more parameter-efficient architectures that provide a better privacy-utility trade-off under data scaling? 3. Contrastive learning remains state-of-the-art for learning features for downstream tasks such as retrieval. What techniques can enable effective DP contrastive learning?

### v2

@@ 4.1 Downstream Tasks
Zero-shot image classification (V-L) is one of the most widely used methodologies for evaluating vision-language models (Radford et al., 2021). A strong zero-shot performance suggests that the image representation aligns well to text. We perform zero-shot classification using the DP-Cap image encoder and text decoder by evaluating the likelihood of captions of the form “this is a photo of a [label]”. We enumerate over different labels and predict the class that has the highest likelihood; see Section A.3.1 for full details.
@@ 4.2 Experimental Setup
We present an overview of the experimental setup; refer to Appendix A for additional details.
Table 2: Linear probing evaluation on downstream classification. Results for DP-NFNet, TAN, AlexNet and SimCLR are obtained from Yu et al. (2023). For ViP (Yu et al., 2023), we train with the same privacy parameters as for DP-Cap on the deduplicated dataset. More details are given in Appendix A.1.
@@
Table 3: Performance of DP-Cap on zero-shot classification and compositional understanding (ARO). CLIP’s zero-shot results are obtained from Radford et al. (2021) (base model). For ARO, see Appendix A.3.2.
@@
We use the ImageNet-1K (Deng et al., 2009a; Russakovsky et al., 2014), CIFAR-10/100 (Krizhevsky et al., 2009), Places-365/205 (Zhou et al., 2014) and iNaturalist-2021 (Van Horn et al., 2021) image classification datasets to assess the performance of learned image representations via full linear probing, few-shot linear probing, and zero-shot prediction. For vision-language tasks, we employ the Visual Genome Attribution (VGA), Visual Genome Relation (VGR), COCO-order (Lin et al., 2015) and Flickr-30k (Plummer et al., 2016) datasets from the ARO benchmark (Yuksekgonul et al., 2022). Finally, we evaluate image captioning using the MS-COCO 2017 (Lin et al., 2015) test set; result is shown in Figure 1(c) and Appendix B.3.
Model and training. We use a transformer architecture (Vaswani et al., 2017) for both the encoder and the decoder of DP-Cap, where the decoder applies causal cross-attention. We train DP-Cap model from scratch, including the embedding layer, with vocabulary size around 32,000. See Section A.1 and Tschannen et al. (2023) for details. For privacy accounting we use Rényi DP composition along with privacy amplification via Poisson subsampling (Mironov et al., 2019), and convert to DP using Balle et al. (2020) through the Opacus library (Yousefpour et al., 2021), targeting $\delta=1/N$ where $N$ represents the number of training samples. We refer to the non-private counterpart of DP-Cap trained on the same dataset as “Cap”.
@@ 4.3 Main Results
Linear probing evaluation (V). We assess the performance of the vision encoder on downstream tasks via linear probing. In Fig 1(a), we compare the performance of DP-Cap and ViP (Yu et al., 2023) on ImageNet-1k few-shot linear probing. DP-Cap significantly improves over ViP, with up to $\times$2.5 better performance across different shots. In addition, we evaluate the full linear probing accuracy of DP-Cap, ViP and other baselines in Table 2. DP-Cap outperforms ViP and other DP models, including TAN (Sander et al., 2023) and DP-NFNet (De et al., 2022), across all tasks. DP-Cap even outperforms non-private AlexNet (Krizhevsky et al., 2012) and except on ImageNet, SimCLR (Chen et al., 2020) (both were trained on ImageNet). We provide additional results for fine-tuning on downstream datasets in Table 10 (App. B), also showing improvements over competing methods.
@@ 4.4 Ablation Studies
We perform ablation studies on the scaling behavior of DP-Cap with respect to the dataset size, privacy budget and model size. In Appendix B, we show additional results on image captioning and on the impact of compute budget.
@@
Impact of the privacy budget $\varepsilon$. We also investigate the performance of DP-Cap under lower privacy budgets ($\varepsilon=1$ and $\varepsilon=2$), employing the same batch size of 1.3 million. The outcomes of these experiments are presented in Table 4. As anticipated, the utility of our model does exhibit a decline with decreasing $\varepsilon$. However, the performance degradation is relatively minor for the learned representation, with 10-shot ImageNet performance decreasing from 31.8% ($\varepsilon=8$) to 19.9% ($\varepsilon=1$). More surprisingly, the performance impact on ARO is nearly negligible. It is noteworthy that both models continue to outperform previous state-of-the-art DP models trained with $\varepsilon=8$ (see Figure 1). This phenomenon can be attributed to the relatively small effective noise resulting from the extreme batch size, which for $\varepsilon=1$ remains five times smaller than that used in (Yu et al., 2023).
Scaling model size. Scaling up the model size is one of the most effective approaches for training better non-private foundation models (Brown et al., 2020; Bommasani et al., 2021; Touvron et al., 2023). However, an intuitive understanding of DP-SGD  [^2] This is because the added noise has $L_{2}$ norm $\approx\sigma C\sqrt{d}/B$, where $d$ is the number of model parameters, whereas the gradient norm is constrained to $C$ regardless of model size. suggests that scaling up model size does not improve utility in DP training since more model parameters will lead to lower signal-to-noise ratio, especially when training large DP models from scratch. Recent work (Li et al., 2022b) demonstrates that during DP fine-tuning, gradients occupy a lower effective dimensional subspace, and larger networks can enhance the performance of DP-SGD. To test this hypothesis in pre-training, we train DP-Cap with different model sizes (Tiny, Small, Base, Large) using the same hyperparameters and evaluate their performance in Table 5,11; see Table 6 for details about different model sizes. We observe consistent improvements when scaling up the model from DP-Cap-Tiny to DP-Cap-Large. Our observation suggests that DP-Cap has strong model scaling behavior even with DP-SGD training.
@@ 5 Discussion and Future Work
Through our study we also identify three open problems in the general direction of DP pre-training of large-scale foundation models that are difficult to handle with existing techniques:
  1. 1.
    Scaling up the batch size to extreme levels is crucial for reducing the effective noise and facilitating model convergence. Is there a fundamental reason why image captioning can support extremely large batch training? When the model does not scale well with batch size, is it due to poor optimization or is it inherent to the model and/or task? Answers to these question can help identify other successful training recipes that further improve the quality of DP-learned representations.
  2. 2.
    Scaling up the model size is typically beneficial in non-private representation learning, but can be ineffective in DP training due to a decrease in gradient SNR. While we have observed a performance improvement from Tiny to Large, we anticipate that deterioration could occur when scaling up to a much larger model size, especially when the signal-to-noise ratio (SNR) undergoes a certain threshold. Does this threshold exist? Can this trade-off be resolved through parameter-efficient architectures?
  3. 3.
    Contrastive learning offers unique advantages compared to other multi-modal learning objectives such as learning image and text representations that are semantically aligned, but is not compatible with standard DP-SGD training. What techniques can enable differentially private contrastive learning?
Acknowledgements
This research was partially supported as a BAIR Open Research Common Project with Meta. We thank Quentin Garrido for his time and insights regarding the significance of vision backbones quality evaluations.
Impact Statement
Our work focuses on developing privacy-preserving techniques for machine learning. Progress in this direction has significant potential for evolving the concept of privacy for AI. The purpose of our work is to push the boundary of what is achievable for representation learning through differentially private mechanisms, towards determining the technical feasibility of applying DP to tackle real-world AI privacy challenges.
Appendix A Implementation Details
A.1 Training Details
DP accounting. We use RDP accounting with subsampling from the Opacus library (Yousefpour et al., 2021). Let $D_{\alpha}$ denote the Rényi divergence of order $\alpha$ (Rényi, 1961), and let
$g_{\alpha}(\sigma,q):=D_{\alpha}{((1-q)\mathcal{N}(0,\sigma^{2})+q\,\mathcal{N}(1,\sigma^{2})\,\|\,\mathcal{N}(0,\sigma^{2}))}.$ (4)
Then, from Mironov et al. (2019), performing $S$ steps of DP-SGD satisfies $(\varepsilon,\delta)$-DP with:
$\varepsilon:=\min_{\alpha}\Big{\{}S\cdot g_{\alpha}(\sigma,q)+\frac{\log(1/\delta)}{\alpha-1}\Big{\}}.$ (5)
The quantity $g_{\alpha}(\sigma,q)$ can be upper bounded mathematically or derived numerically: we use the Opacus (Yousefpour et al., 2021) library for accounting in our work.
Regarding the DP guarantee, $\varepsilon$-DP bounds the amount of information extracted from each training sample by $\varepsilon$. Notably, for DP-Cap, each sample is made of {image + caption}, while ViP utilizes {image} only. Consequently, DP-Cap inherently offers an equivalent or better privacy guarantee for each image. One way to see it is to note that DP provides protection against membership inference attacks (Shokri et al., 2017). Suppose $\varepsilon$-DP upper bounds the success rate of a membership inference attack (when given the image-text pair) against DP-Cap as $\leq p$. Then the MIA success rate when given only the image can be at most $p$ since the attacker has strictly less information. This is exactly the upper bound for the success rate of a membership inference attack against ViP. In other words, any attacker that can attack the image+caption model (such as DP-Cap) can also attack the image-only model (such as ViP).
On the other hand, since the {image+caption} models utilize the caption, the privacy leakage from the text part of the image-caption pair is non-zero for $\varepsilon>0$. It is worth noting that in our set up since we use DP, we protect the captions with the same $\varepsilon$-DP guarantee. Thus, the privacy protection for DP-Cap is neither strictly stronger nor strictly weaker than that for ViP, so the two privacy notions are not directly comparable.
Model details and task description. We utilize a transformer architecture (Vaswani et al., 2017) DP-Cap. This captioning model uses a text decoder that generates captions in an auto-regressive manner, utilizing a full attention mechanism on the vision encoder’s output, as well as causal attention on the text. This architecture is closely aligned with the Cap architecture introduced in (Tschannen et al., 2023). See Table 6 for details about the transformer architecture for different sizes. All results utilize the base model with the exception of the comparison in Table 6.
Hyperparameters. Our choice of gradient clipping factor is $C=1$, as we did not observe any performance improvement with other values. We always use AdamW (Loshchilov & Hutter, 2018) for training. We use a learning rate of $5.12\times 10^{-4}$. The learning rate is kept constant across batch sizes for TAN simulations and for the performance comparison in Figure 3 as the effective noise is kept constant in these cases (Sander et al., 2023). We use a maximum length of $40$ tokens to process the LAION captions. We use a linear schedule, with $40\%$ of warm-up iterations, and 2×2$\times$2 × the entire training as decay horizon. As opposed to what was previously observed (De et al., 2022; Sander et al., 2023), the learning rate schedule played an important role for us with DP-SGD training. We use a weight decay of $0.05$. These choices come from hyperparameter search using TAN simulation with our base model. Following the standard practice (Berrada et al., 2023; De et al., 2022; Li et al., 2021; Yu et al., 2023; Sander et al., 2023), we do not count hyperparameter search within our privacy budget. (Liu & Talwar, 2019) have shown that hyperparameter search might not incur observable privacy loss.
Pre-training DP-Cap on the synthetic dataset. Compared to the encoder and decoder architecture design used in masked autoencoders (MAE) (He et al., 2022), the two main differences of the image captioning model used in this paper are: (1) The output of the encoder is fed into the decoder via cross-attention (Vaswani et al., 2017) in Cap; and (2) The self-attention used in the Cap decoder is causal self-attention. Similar to Yu et al. (2023), we apply the synthetic image dataset, Shaders21K (Baradad et al., 2022), to pre-train the DP-Cap model via MAE-based training. We follow most of the training setups used in ViP synthetic pre-training (Yu et al., 2023), except that we feed the output of the encoder to the decoder via cross-attention. The training loss of the synthetic pre-training in this paper is still the reconstruction loss used in MAE (He et al., 2022), and we did not leverage real-world text data for pre-training. After the model is pre-trained on Shaders21K, we change the self-attention to causal self-attention in the decoder, and replace the final layer (for pixel-wise reconstruction) of the decoder with the (randomly initialized) decoding layer for next word prediction. After making these modifications, we apply DP-SGD to pre-train our DP-Cap model with standard image captioning training objectives (see Section 3.1).
Table 6: Details of transformer backbone variants used in DP-Cap.
Model Encoder depth Encoder width Decoder depth Decoder width # parameters (encoder & decoder)
DP-Cap-Tiny 12 384 6 384 59M
DP-Cap-Small 12 576 6 576 115M
DP-Cap-Base 12 768 6 768 190M
DP-Cap-Large 24 1024 6 768 407M
Pre-training ViP. To conduct a comparison with training on an identical datasets, we follow the methodology outlined in (Yu et al., 2023) to train with DP-SGD a MAE-based model, but with a change in the training data from LAION-233M to Dedup-LAION-223M, and use the same encoder’s synthetic initialization as for DP-Cap. We further examine the linear probing performance on ImageNet and observe a 2% between the original model and the one trained on the deduplicated dataset. In addition, to corroborate the observation made in Figure 3, which suggests that the MAE-based method struggles to effectively harness massive batch sizes for achieving low effective noise in DP-SGD, we also train ViP models with larger batches, up to using the exact privacy parameters employed for DP-Cap (under $\varepsilon=8$) with a notably large batch size of 1.3 million, and showcase the results in Table 7. For full linear probing, we observe only a small improvement over the original ViP model that was trained with batch size 98k. The success of DP-Cap is not solely attributed to its appropriate privacy parameters but is also a consequence of its remarkable ability to leverage the small effective noised induced by extremely large batch sizes.
A.2 Computation cost
Mixed Precision Package & Ghost Norm
DP-SGD introduces additional computational overhead compared to non-private training, primarily due to the computation of per-sample gradient norms. By employing the ghost norm technique of Li et al. (2021), we have successfully reduced the computational cost by up to one third with the Large Model (see Figure 4) compared to using functorch. The torch.amp package offers convenient methods for mixed precision, significantly speeding up operations like linear layers. However, it often leads to NaNs due to low precision handling of extreme values. While one can skip a step that led to NaNs in normal training, combining AMP with Ghost Norm is more complex. Ghost Norm requires two backward passes. In the first pass, per-sample gradient norms are computed. If one gradient norm is NaN, it contaminates the entire batch, leading to a NaN in the second backward pass. This issue is particularly prevalent in our setting with a batch size of $1.3$M, as even a minuscule proportion of computations leading to NaNs can cause problems. To address this, we propose two solutions:
  • Loss Scaler: We employ a similar trick to the standard use of AMP to reduce the number of NaNs. This involves dynamically upscaling and downscaling the loss with torch.cuda.amp.GradScaler. The same factor is used before the first and the second backward, and is updated based on the outputs of the second backward only.
  • Clipping to 0: If any per-sample gradient norm computation results in a NaN value after the first backward, we set its clipping coefficient (the multiplicative coefficient in front of the corresponding per-sample loss for the second backward, as detailed in Li et al. (2021)) to 0 for the second backward. In this case, we do not update the loss scaling factor.
It’s worth noting that the second solution is entirely valid for DP: instead of clipping the per-sample gradient to a norm $C$, it clips it to 0 in cases where computation results in a NaN value. This approach effectively mitigates the issue of NaN contamination in large batches. Overall, We have successfully reduced the computational cost by a factor 5 for the Large Model compared to functorch.
TAN simulation
Crucially, to achieve a favorable privacy-utility trade-off, DP-SGD necessitates training with massive batches over a substantial number of steps to achieve a good privacy-utility trade-off, as elaborated in Section 3.2. All our hyperparameter search were performed using the TAN simulation (Sander et al., 2023) for one epoch on our Dedup-LAION-233M. For our $\varepsilon=8$ models, we limited training to 32 epochs, a process that took 5 days utilizing 128 V100 GPUs for the Base model.
While we have tried to reduced it as much as possible, training DP-Cap imposed a considerable energy consumption, resulting in elevated CO2 emissions. Our intention in releasing these models is to contribute to the mitigation of future carbon emissions, as the training has already been completed.
Table 7: Set-ups for our training for ViP, MAE (He et al., 2022) and DP-Cap: ImageNet-1k linear probing.
Model pretraining data ($B$, $\sigma$, $S$) ImageNet-1K
1-shot 2-shot 5-shot 10-shot full
ViP (Yu et al., 2023) LAION-233M (98k, 0.48, 6000) 2.5% 4.1% 8.5% 14.2% 55.7%
ViP Dedup-LAION-233M (98k, 0.474, 5708) 2.3% 3.9% 8.0% 13.6% 53.4%
ViP Dedup-LAION-233M (200k, 0.513, 5708) 2.5% 4.1% 8.3% 14.1% 54.0%
ViP Dedup-LAION-233M (400k, 0.564, 5708) 2.5% 4.2% 8.7% 14.7% 55.2%
ViP Dedup-LAION-233M (1.3M, 0.728, 5708) 2.7% 4.6% 9.4% 15.7% 56.5%
MAE (Non private) Dedup-LAION-233M (40960, 0, 40000) 3.4% 5.8% 11.8% 19.5% 62.5%
DP-Cap Dedup-LAION-233M (98k, 0.474, 5708) 4.2% 6.9% 11.4% 17.2% 50.6%
DP-Cap Dedup-LAION-233M (200k, 0.513, 5708) 5.6% 8.9% 14.5% 20.2% 54.2%
DP-Cap Dedup-LAION-233M (400k, 0.564, 5708) 7.6% 11.5% 18.5% 25.3% 59.1%
DP-Cap Dedup-LAION-233M (1.3M, 0.728, 5708) 10.3% 15.6% 24.2% 31.8% 63.4%
A.3 Evaluation Details
A.3.1 Details about Zero-shot Image Classification
While methods employing contrastive learning, such as CLIP, excel in this task, captioning methods exhibit comparatively lower performance, and with greater computational demands during evaluation. To evaluate a captioning model’s zero-shot performance, we employ two distinct strategies:
  • Tree-based search: We initiate caption generation with a prompt like “this is a photo of a,” and greedily select the most likely next token among those that lead to valid completions within the true label set. The process continues until an End of Sentence (EOS) token is reached. For instance, if there are only two labels starting with “car”: “car [EOS]” and “carpet [EOS]”, and the initial predicted token is “car”. Then the text decoder will predict the next token among “[EOS]” and “pet”. If, among these two, “[EOS]” is chosen, and “car [EOS]” corresponds to the true label, then the zero-shot prediction is deemed correct.
  • Loss-based classification: We assess, for each image, the probability of various captions that begin with “this is a photo of a […]” where “[…]” is substituted with all feasible labels. Subsequently, we select the label that yields the most probable caption.
The “loss-based classification” comes with significantly higher computation costs as all the different captions have to be evaluated for each image (there representations is conditional to the image). For ImageNet, it implies 1000 forwards through the decoder for each image. We thus employ the tree-based search for presenting our findings in Table 3, although its greedy character with no backtracking is not optimal. Surprisingly, our preliminary experiments suggest the tree-based search gives comparable results.
A.3.2 Details about ARO Evaluation
We adhered to the protocol and code base established in (Yuksekgonul et al., 2022) for re-evaluating CLIP’s performance, and we observe slightly different results (see Table 8). For our captioning models, our approach involved computing the cross-entropy loss for all possible captions associated with each image and subsequently selecting the one with the lowest loss.
Table 8: Compositional understanding (ARO): Results for CLIP (base) in Yuksekgonul et al. (2022) compared to our evaluation.
Model ARO
VGR VGA COCO Flickr
CLIP (eval from Yuksekgonul et al. (2022)) 59% 63% 46% 60%
CLIP (our eval) 62.4% 62.9% 47.8% 58.0%
A.3.3 Details about Linear Probing and Fine-tuning Evaluation.
Few-shot linear probing is accomplished using the Cyanure library (Mairal, 2019). We use the same hyper parameters as in Assran et al. (2022). We adapted the MAE (He et al., 2022) code base for full linear probing, and we use the same hyperparameters as in (Yu et al., 2023) (extract 12 layers of the image encoder, LARS optimizer (You et al., 2017) with base learning rate of 0.1, no weight decay and batch size of 16384).
Appendix B Additional Results
B.1 Additional experiments
Impact of the initialization (V). Our synthetic initialization for DP-Cap achieves less favorable results than the one from ViP reaches 50% (Yu et al., 2023); for instance, for full linear probing on ImageNet, they achieve 44% (Figure 2) and 50% respectively. However we have demonstrated that training with DP on top of synthetic initialization leads to significantly better results for DP-Cap compared to ViP for all the metrics; see Table 2, Table 10 and Figure 1. We observe that this superiority also appears when the models are trained from random initialization: as shown in Table 9, the improvement over ViP is even larger when training without synthetic initialization.
Table 9: Training from random initialization: Superiority of DP-Cap over ViP, both trained from random initialization.
Model ImageNet-1K
1-shot 2-shot 10-shot full
ViP ($\varepsilon=8$) 0.1% 1.7% 6.1% 23.9%
DP-Cap ($\varepsilon=8$) 5.6% 8.5% 18.8% 47.0%
Fine-tuning (V). In Table 10, we present DP-Cap’s performance in fine-tuning for few-shot evaluation. In contrast to the linear probing results shown in Table 2, the network is completely unfrozen. Therefore, we assess DP-Cap’s capabilities primarily as a network initialization. Similarly to the linear probing results, we note a significant improvement in all metrics compared to previous DP vision backbones. Note that, similarly to linear probing comparison in Figure 1, we compare to non-private model performance which provides information about the performance gap between private models and non-private models. For fair comparison, we evaluate on the same same datasets than Yu et al. (2023).
Table 10: Fine-tuning evaluation on few-shot downstream classification.
Model Aircraft Caltech-101 CIFAR-100
10-shot 20-shot 30-shot 5-shot 10-shot 30-shot 5-shot 10-shot 30-shot
AlexNet 23.3% 34.4% 41.4% 64.7% 73.6% 81.4% 29.7% 36.3% 49.3%
SimCLR 38.8% 56.9% 64.9% 81.7% 89.1% 94.5% 49.9% 60.2% 71.8%
TAN 22.8% 37.9% 46.0% 49.3% 66.4% 77.9% 21.3% 27.8% 42.4%
ViP 31.6% 53.1% 64.3% 68.1% 79.0% 88.9% 30.7% 41.0% 57.5%
DP-Cap 37.5% 57.9% 66.7% 70.3% 81.3% 90.0% 36.3% 46.3% 62.1%
Table 11: Ablation studies on the effect of model size for zero-shot prediction.
Model Config # parameters DP? Zero-shot
ImageNet-1k CIFAR10 CIFAR100
Random Chance - - - 0.1% 10% 1%
Cap Base 86.6M 25.2% 90.0% 37.4%
DP-Cap Tiny 22.0M 5.0% 46.5% 11.1%
DP-Cap Small 49.0M 6.9% 53.6% 17.1%
DP-Cap Base 86.6M 7.8% 54.4% 16.4%
DP-Cap Large 407.3M 9.2% 62.1% 24.0%
Captioning task (V-L). We evaluate the image captioning performance of DP-Cap in comparison to non-private Cap. In Fig. 1(c), we present some (randomly chosen) captions generated by DP-Cap; more examples for DP-Cap and Cap can be found in Appendix B.3. Qualitatively, DP-Cap seems to generate reasonably good captions, similar to the ones generated by Cap. We also compare the two models quantitatively using the CIDEr metric (Vedantam et al., 2015) to evaluate the generated captions on the MS-COCO test set, and the results are summarized in the last column of Table 3. As DP-Cap and Cap are only trained on noisy captions from LAION, the CIDEr metric on MS-COCO is relatively low for both models. Moreover, despite the similar performance between DP-Cap and Cap on ARO, the gap is much more significant for the captioning evaluation. Given these results, it is plausible that even though DP-Cap attains remarkably compositional understanding capabilities, its ability to generate text is still limited.
We also fine-tune Cap and DP-Cap’s decoders (while freezing the encoder) for one epoch on the MS-COCO train set, and assess the improvement in CIDEr scores in Table 12 to showcase the quality of the image representations and decoder initialization from the pre-training stage. The captions in Figure 1 and Appendix B.3 are generated using models that were not trained on MS-COCO.
Table 12: Captioning evaluation on the MS-COCO test set of Cap and DP-Cap. For “fine-tuned”, the model’s decoder is fine-tuned for one epoch on the MS-COCO train set (with the image encoder frozen).
Model CIDEr score
original fine-tuned
Cap 29.9 79.2
DP-Cap 15.7 51.3
What can we do with more compute budget? We restricted training the DP-Cap model for a compute budget of 32 epochs on the Dedup-LAION-233M dataset for each of our models with $\varepsilon=8$. To fit the privacy budget while utilizing a batch size of 1.3 million and training for 32 epochs, RDP analysis yields $\sigma=0.728$. However, we anticipate that further increasing the compute budget can yield even better models up to a certain limit: With the same $\varepsilon$ and batch size, doubling the compute to 64 epochs only necessitates a 12% increase in $\sigma$. This increase enables twice as many steps to be performed with only a marginal increase in effective noise, potentially allowing the model to converge to a better solution.
In the absence of necessary compute for running this experiment, we partially validate this hypothesis through the Total Amount of Noise (TAN) simulation, training for the same number of gradient steps and with the same SNR per step, but using a $\times$32 smaller batch size and $\times$32 smaller $\sigma$ to simulate at $\times$32 lower compute. Our results in Table 13 indicate a significant performance improvement of 5% in 10-shot accuracy on ImageNet (compared to a similar simulation of the 32 epochs training). However, increasing the budget further to 128 epochs does not seem to enhance performance compared to the 64 epoch counterpart. Intuitively, the lower gradient SNR and larger number of gradient steps have opposite effects on optimization, and pushing past the “sweet spot” of training for 64 epochs at $\sigma=0.81$ results in noisy steps that are unproductive for model convergence. To surpass the performance of the 64-epoch, 1.3-million batch size DP-Cap model, training with an even larger batch size appears necessary. We emphasize again that this result is derived through TAN simulation, and actual, compute-intensive training is required to fully validate this assertion.
Table 13: TAN simulation of the impact of the compute budget on the performance at fixed $B$.
$\sigma$
0.81 0.95
Epochs 64 ($\times$2) 128 ($\times$4)
Effective noise $\sigma/B$$\times$1.12 $\times$1.32
Predicted Final loss --0.2 --0.2
Predicted 10-shot ImageNet $+$5% $+$5%
B.2 More on the Impact of dataset size and privacy parameters
Dataset size. We emphasize here (again) the importance of having enough training data to achieve a good privacy-utility trade-off with DP-SGD. As depicted in Figure 5, increasing the number of training samples $N$ while keeping the same number of equivalent DP-SGD steps (i.e., keeping batch size $B$, noise $\sigma$, and number of update steps $S$ constant) considerably reduces the privacy budget $\varepsilon$. Equivalently, having more data allows for an increase in the number of equivalent DP-SGD steps at fixed $\varepsilon$. Similar observations were also made by Tramer & Boneh (2020); McMahan et al. (2017). The abundance of pre-training data available for training foundation models thus proves highly compatible with DP requirements.
Figure 5: At fixed (B, $\sigma$, S), $\varepsilon$ drastically reduces with the dataset size.
Figure 6: All points correspond to $\varepsilon=8$ for a dataset of size $N=233$M. At fixed $\varepsilon$ and $\sigma$, the number of epochs decreases as the batch size increases.
Batch size and $\sigma$. We wish to underscore the influence of batch size and $\sigma$ on both the computational budget and model performance. As highlighted in Section 3.2, for a given target $\varepsilon$, elevating $\sigma$ beyond 0.5 allows training for significantly more steps. In Figure 6, the blue, orange and green lines show the batch size ($B$) vs. compute trade-off ($E$) at a given $\sigma$. The lines are monotonically decreasing with $B$, signifying that the number of epochs $E$ decreases when increasing $B$. When maintaining a fixed privacy budget $\varepsilon=8$, even a marginal increase in $\sigma$ from $0.48$ to $0.728$ (from blue to orange) translates to a remarkable increase ranging from 100 (for small batch sizes) to 100,000 (for very large batch sizes) times more gradient steps. Thus it is favorable to increase $\sigma$ and $B$ at the same time for better model convergence.
Meanwhile, doing so also incurs a higher computational cost: Under a 32-epoch budget on Dedup-LAION-233M with a batch size of $1.3$ million, we had to cut the red curve in Figure 6, with $\sigma=0.728$. As outlined in Section 4.4, with twice this budget, we could have raised $\sigma$ to 0.81 (green curve), with simulations indicating that this would have substantially improved performance. Additionally, Section 3.2 underscores that increasing the batch size is pivotal for achieving a high SNR while maintaining reasonable privacy guarantees. It is also crucial to note that at fixed $\varepsilon$, the compute budget is inversely proportional to the batch size. Therefore, increasing the batch size is beneficial for both SNR and computational efficiency. However, an excessively large batch size leads to fewer epochs and consequently a very limited number of training steps, which is detrimental to the training process (in addition to the difficulties of large batch training). For optimal privacy-utility-compute trade-off, a balance must be struck between computational resources, feasible batch size, and a reasonable number of training steps.
B.3 Image Caption Examples
In Figures 7 and 8, we show images from the MS-COCO 2017 test set and their corresponding captions generated by human annotator, Cap, and DP-Cap. Images in Figure 7 are selected randomly, whereas images in Figure 8 are randomly selected from the top 10% CIDEr score examples for DP-Cap. Qualitatively, the human-generated captions are more precise, whereas the captions generated by Cap and DP-Cap are more generic and sometimes contain factual errors. This is to be expected since Cap and DP-Cap are trained on LAION with much noisier text description and were not fine-tuned on MS-COCO. Nevertheless, DP-Cap still generates grammatically correct and (mostly) semantically coherent captions for unseen images.
Figure 7: Captions of randomly selected images from the MS-COCO 2017 test set.
Figure 8: Captions of images (randomly picked among the top 10% CIDEr score of DP-Cap) from the MS-COCO 2017 test set.